what are the dimensions of a parallelogram with the area of 135 with the height = x-3 and base = x+3
Answer by stanbon(75887)

what are the dimensions of a parallelogram with the area of 135 with the height = x-3 and base = x+3
-----
base*height = 135
(x+3)(x-3) = 135
-----
x^2 - 9 = 135
x^2 = 144
x = 12
----
base = x+3 = 15
height = x-3 = 9
---------------------
Cheers,
Stan H.
====================